OpenStudy (anonymous):

Will give medals!! The perimeter of a rectangle is 68 ft. The ratio of it's length to its width is 9 : 8. What are the dimensions of the rectangle?

OpenStudy (anonymous):

18 ft by 16 ft is the answer.

OpenStudy (anonymous):

oh your welcome btw

OpenStudy (chrisisdope1):

Its not 18 by 16 it 18 and 16

OpenStudy (anonymous):

oh sry lol

OpenStudy (anonymous):

I had a brainfart for a sec. :)

OpenStudy (midhun.madhu1987):

Perimeter = 68 Length:breadth = 9:8 So, let Length = 9x and breadth = 8x Perimeter of the rectangle , P = 2(9x + 8x) 68 = 2(9x + 8x)

OpenStudy (anonymous):

Im in 7th grade and I knew that ha. lol